rolly is 9 years younger than greg. in seven years, he will be two years more than one-half as old as greg. find their present ages.
Answer by ptaylor(2198)

Let x=Greg's present age
Then x-9=Rolly's presentage
In 7 years, Greg's age will be x+7
In 7 years, Rolly's age will be (x-9)+7 or x-2
Now we are told that:
(x-2)=1/2(x+7)+2 multiply each term by 2
2(x-2)=(x+7)+4 get rid of parens (distributive law)
2x-4=x+7+4 subtract x from and add 4 to both sides
2x-4+4-x=x-x+7+4+4 collect like terms
x=15-------------------------Greg's present age
x-9=15-9=6-----------------------Rolly's present age
CK
6+7=1/2(15+7)+2
13=11+2
13=13

Hope this helps---ptaylor